# Through the Dual-Prism: A Spectral Perspective on Graph Data Augmentation for Graph Classification

### v1

@@ Appendix E More Discussions
Rationale for Choosing Graph Laplacian Decomposition. In our methodology, we chose to decompose the graph Laplacian $L$ (where $L=D-A$) to do the perturbation and then reconstruct the graph. In terms of implementation, an alternative method can be directly decomposing $A$ and perturbing its smaller eigenvalues. However, our motivation for this work is more on the inherent properties of graphs, and $L$ offers a more nuanced reflection of these properties compared to $A$ Lutzeyer & Walden (2017). For future scenarios involving more complex disturbances to eigenvalues, leveraging the eigenvalues of $L$ would be a more appropriate approach. In addition, our decision to decompose $L$ also follows general spectral graph convolution methodologies Kipf & Welling (2016a).

### v3

@@ Acknowledgements
Rationale for Choosing Graph Laplacian Decomposition. In our methodology, we chose to decompose the graph Laplacian $L$ (where $L=D-A$) to do the perturbation and then reconstruct the graph. In terms of implementation, an alternative method can be directly decomposing $A$ and perturbing its smaller eigenvalues. However, our motivation for this work is more on the inherent properties of graphs, and $L$ offers a more nuanced reflection of these properties compared to $A$ (Lutzeyer and Walden 2017). For future scenarios involving more complex disturbances to eigenvalues, leveraging the eigenvalues of $L$ would be a more appropriate approach. In addition, our decision to decompose $L$ also follows general spectral graph convolution methodologies (Kipf and Welling 2016a).
Loss-pass Filtering of GNNs vs. DP Augmentation. We then discuss the relationships and differences between the loss-pass filtering of GNNs, caused by the model’s neighborhood aggregation processes (Nt and Maehara 2019; Kipf and Welling 2017), and our proposed methods.
  • Relationship. The low-pass effect of GNNs exactly aligns with our motivation - increasing the diversity of data (i.e. altering the high-frequency part) without misguiding the classification (i.e. preserving the low-frequency part). In other words, because DP augmentation enhances the training-data diversity by introducing randomness to the high-frequency spectrum, the low-pass effect of GNNs preserves the core attributes necessary for correct classification, ensuring label consistency across original and augmented data. By using this natural smoothing capability, our method introduces label-consistent variations to the data. This strategy boosts model robustness and generalization without compromising classification accuracy, affirming the efficacy and theoretical integrity of our method.
  • Difference. From the signal processing perspective, our method is more like introducing random perturbations to high-frequency components rather than acting as another low-pass filter. For better intuitive understanding, take the image processing as an example, a low-pass filter achieves image denoising, while random perturbation of high-frequency noise sharpens images. These two effects are different and even have a certain degree of opposition. Due to such difference, these two operations can be used jointly during the image processing - in our work, we first use augmentation to randomly perturb the high-frequency part, and then pass it through the low pass filter of GNNs.
Possible reasons of why DP-Noise outperforms DP-Mask. DP-Noise applies subtle, controlled noise to high-frequency eigenvalues, preserving essential low-frequency components while introducing diversity through minor spectral changes. This balance allows DP-Noise to produce augmented graphs that are both diverse and stable, enhancing model generalization and robustness. Conversely, DP-Mask removes selected high-frequency eigenvalues outright, creating more pronounced structural shifts. While effective in diversifying, this approach can introduce instability by over-modifying spectral components, impacting performance consistency.